OpenStudy (anonymous):

convert 13/25 into a decimal

OpenStudy (tkhunny):

1) Get your calculator to divide it for you. 2) If you're quick, you can do the division by yourself. 3) Multiply numerator and denominator by 4.

OpenStudy (anonymous):

im slow :D

OpenStudy (anonymous):

but ok

OpenStudy (tkhunny):

I recommend #3, but you do have to see it.

OpenStudy (anonymous):

13/25*4/4=52/100=0.52

OpenStudy (anonymous):

@.@

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw